THE CIRCLE TRANSFORM ON TREES

ENRICO CASADIO TARABUSI, SIMON G. GINDIKIN, AND MASSIMO A. PICARDELLO

Abstract. We consider the overdetermined problem of integral geometry on trees given by the transform that integrates functions on a tree over circles, and exhibit difference equations that describe the range. We then show how this problem modifies if we restrict the transform to some natural subcomplex of the complex of circles, proving inversion formulas and characterizing ranges.

1. Introduction

There is a strong analogy between the geometry and the harmonic analysis of trees and those of real hyperbolic spaces [BCP], [Ca]. This analogy is better understood by regarding trees as p-adic symmetric spaces of rank one [S], [M]. It is natural to interpret inversion formulas for the horospherical and geodesic Radon transforms on trees as discrete versions of corresponding formulas for hyperbolic spaces [BCCP], [CCP], and vice versa.

Overdetermined problems play an important role in integral geometry. The classical example is F. John’s transform [J], which maps functions on the three-dimensional Euclidean space to their integrals over lines. Thus a function of three variables is transformed into a function of four variables. As a result, the range consists only of functions that satisfy the ultrahyperbolic equation. Correspondingly, in inversion formulas it is natural to use the values of the transformed function not on all lines, but only on those in some three-dimensional subfamily (called complex).

Another classical overdetermined problem is given by the circle transform in the plane (or, more generally, the sphere transform). This maps functions of two variables to functions of three variables, and we get again the situation described above: a description of the range by a differential equation, and inversion using two-dimensional families of circles.

Perhaps for the first time, we consider an overdetermined problem on trees, namely the discrete analogue of the circle transform: Given a function f(v) on the vertices of a tree, we take f̂(v, n) to be its “integral” on the circle of center v and radius n. To work in greater generality, we shall in fact also allow the center to be an edge of the tree. In some sense, f̂ depends on one extra variable with respect to f, and as a result it must satisfy a suitable equation. Indeed we obtain in Section 3 some difference equations that describe the range of the transform f ↦ f̂.

1991 Mathematics Subject Classification. Primary 44A12; Secondary 05C05, 43A85.

Key words and phrases. Radon transform, trees, circles.

Research partially supported by US National Science Foundation grant DMS-0070816 and the Mathematical Sciences Research Institute.

To eliminate the overdetermination, as in the continuous case we restrict the transform to some families of circles that, in some sense, have codimension one; we choose those that are the analogues of some so-called admissible families of circles in the plane for which there are good inversion formulas. As a basic example, in Section 4 we restrict the circle transform to the subcomplex of circles passing through a given fixed vertex, proving an inversion and a range characterization for ℓ^{1}. We finally consider other subcomplexes for which, by analogy with the continuous case, we can expect good inversion formulas.

2. Preliminaries and notation

Let T be a tree, that is, a connected, undirected, locally finite graph without loops and self-loops. The set of vertices of T will be denoted by V, the set of edges by E, and both will be endowed with the counting measure for integration purposes. Two vertices u, v are neighbors, and we write u ∼ v, if they are connected by an edge, which in this case is unique, is denoted by uv and identified with the set {u, v}. The degree of a vertex v is the number q_{v} + 1 of its neighbors; if this is a constant q + 1, then the tree is homogeneous. The interior of a subset K of V is the set of its vertices whose neighbors are all in K. Two edges uv, wz are neighbors, or uv ∼ wz, if they share exactly one vertex. An edge uv has exactly q_{u} + q_{v} neighboring edges. The distance between two vertices u, w is the length n of the unique injective path [u, w] = (u = v_{0}, v_{1}, …, v_{n} = w) from u to w for which v_{j−1} ∼ v_{j} for every j = 1, …, n. Instead, the distance between an edge uv and a vertex w is d(uv, w) = min(d(u, w), d(v, w)); it vanishes if and only if w ∈ uv. A subset K of V is connected (or convex) if it contains the path between any two of its vertices. For instance, a geodesic, that is, a doubly infinite path in V, is a connected subset of V.

Let N be the set of non-negative integers. A circle in V is the set

C_{x,n} = {u ∈ V : d(x, u) = n} for x ∈ V ∪ E and n ∈ N,

where x is its center and n its radius. Although all circles are subsets of V, the above definition thus encompasses those whose centers are vertices (henceforth v-circles) as well as those whose centers are edges (e-circles). The space C of circles is thus in one-to-one correspondence with (V ∪ E) × N. The circle transform of a function f on V is the function f̂ on C given by

f̂(x, n) = ∑_{u∈C_{x,n}} f(u) for (x, n) ∈ (V ∪ E) × N.

A circle C_{x,n} is externally, respectively internally tangent to a subset K of V if the minimum, respectively maximum distance of x from points of K is n.

We shall assume in Section 4 that a reference vertex o ∈ V is fixed. The length of x ∈ V ∪E is |x| = d(x, o). A vertex v is even or odd if so is its length. Correspondingly, V_{even}, V_{odd} are the subsets of even, respectively odd vertices of V. If 0 ≤ n ≤ |v|, denote by v_{n} the n-th vertex of the path from o to v, so that |v_{n}| = n, and the whole path is

[o, v] = (o = v_{0}, v_{1}, …, v_{|v|} = v).

The vertices v_{0}, …, v_{|v|} are the predecessors of v. In particular, if v ≠ o then v_{|v|−1}, i.e., the unique predecessor of v that is at the same time its neighbor, is also denoted by v^{−}. Setting V′ = V \ {o}, the map μ: V′ → E given by

μ(v) = vv^{−} for v ∈ V′

is a bijection. The set of vertices of which u is a predecessor is

S_{u} = {v ∈ V : |v| ≥ |u|, v_{|u|} = u} for u ∈ V,

and the subset of those at distance n from u is

S_{u,n} = S_{u} ∩ C_{u,n} = C_{o,|u|+n} ∩ C_{u,n} for u ∈ V and n ∈ N.

If x ∈ E set S_{x} = S_{μ^{−1}(x)} and S_{x,n} = S_{μ^{−1}(x),n} for n ∈ N (all are sets of vertices). The join of two vertices v, w is their longest common predecessor.

In order to give a more suggestive expression of the range-characterizing equations we need some notation for finite differences. The forward and backward difference operators ∂^{+}, ∂^{−} act on a function h on N as

∂^{+}h(n) = h(n + 1) − h(n),
∂^{−}h(n) = h(n) − h(n − 1),

where we henceforth set h(−1) = 0. Also, let

∂ = \frac{∂^{+} + ∂^{−}}{2} ,
∂^{(2)} = ∂^{+} − ∂^{−} = ∂^{+}∂^{−} = ∂^{−}∂^{+}

be the symmetric two-step difference, respectively second order difference operators. By i we shall denote the identity operator (i.e., ih(n) = h(n)).

The Laplace operator on V being given by

Δf(v) = \frac{1}{q_{v} + 1} ∑_{u∼v} f(u) − f(v),

then for any real c > 0 the discrete-time wave operator, acting on functions on V ×N, can be defined as

□_{c} = Δ − c∂^{(2)}.

The same expression provides a wave operator on functions on E × N in terms of the Laplace operator on E, which, if T is homogeneous of degree q + 1, is given by

Δf(uv) = \frac{1}{2q} ∑_{wz∼uv} f(wz) − f(uv)

(every edge has 2q neighboring edges).

The backward difference operator D^{−} on V takes a function f to the function Df given by

D^{−}f(v) = f(v) − f(v^{−}) for every v ∈ V,

where we set f(o^{−}) = 0. The inverse of D is given by

(2.1) f(v) = ∑_{w∈[o,v]} D^{−}f(w) for every v ∈ V.

Instead, the forward difference operator on V can be defined by

D^{+}f(v) = ∑_{u^{−}=v} f(u) − f(v) for every v ∈ V.

Lemma 2.1. The operator D^{+} maps ℓ^{1}(V) into itself with operator norm not greater than 2, and it can be inverted, for f ∈ ℓ^{1}(V), by

f(v) = − ∑_{u∈S_{v}} D^{+}f(u) for every v ∈ V,

where the series converges absolutely.

Proof. The assertion about the norm follows from the uniqueness of the immediate predecessor u^{−1} of each given u ≠ o. The series converges absolutely because D^{+}f ∈ ℓ^{1}(V). The inversion formula is obtained by summing for n ∈ N the relation

∑_{u∈S_{v,n}} D^{+}f(u) = ∑_{u∈S_{v,n+1}} f(u) − ∑_{u∈S_{v,n}} f(u) for every v ∈ V and n ∈ N.

We shall need in Section 4 to extend D^{+}, D^{−} to V ∪ E in a somewhat hybrid fashion as

D^{+}f(vv^{−}) = f(v) − f(vv^{−})  for every v ∈ V′,
D^{−}f(vv^{−}) = f(vv^{−}) − f(v^{−}) for every v ∈ V′.

3. The full circle transform

We can now characterize the range of the circle transform on the space C of all circles.

Theorem 3.1. A function ϕ on (V ∪ E) × N is in the range of the circle transform if and only if the following two equations hold:

(3.1) ∑_{u∼v} ϕ(u, n) = ϕ(v, n + 1) + ϕ(v, n − 1) · \begin{cases}anything & if v ∈ V and n = 0, \\ (q_{v} + 1) & if v ∈ V and n = 1, \\ q_{v} & if v ∈ V and n > 1, \end{cases}

ϕ(u, n) + ϕ(v, n) = ϕ(uv, n) + ϕ(uv, n − 1) for uv ∈ E and n ∈ N (3.2)

(with ϕ( · , −1) ≡ 0). Equation (3.1) alone characterizes the range on v-circles, and can be rewritten as Z_{V} ϕ = 0 where

Z_{V} = □_{1/2} + \frac{q_{v} − 1}{q_{v} + 1}∂ + \frac{∂^{−} − i}{q_{v} + 1} · \begin{cases}anything & on V × \{0\}, \\ 1 & on V × \{1\}, \\ 0 & on V × (N \ \{0, 1\}). \end{cases}

If T is a homogeneous tree of degree q + 1, equation (3.2) can be replaced by

(3.2′) \{ϕ(u, 0) + ϕ(v, 0) = ϕ(uv, 0) (this is (3.2) for n = 0), \begin{aligned}∑_{wz∼uv} ϕ(wz, n) & = ϕ(uv, n + 1) + qϕ(uv, n − 1) \\ & + ϕ(uv, n) · \begin{cases}q & if n = 0, \\ (q − 1) & if n > 0, \end{cases} \end{aligned}

for every uv ∈ E.

The second equation of (3.2′), which only involves values at e-circles, can be rewritten as Z_{E}ϕ = 0, where

Z_{E} = □_{(q+1)/4q} + \frac{q − 1}{2q}∂ − \begin{cases}\frac{i}{2q} & on E × \{0\}, \\ 0 & on E × (N \ \{0\}); \end{cases}

the first two summands of this right-hand side can also be written as

□_{1/2} + \frac{q − 1}{2q}∂^{+}.

Proof. Let f be a function on V, and let ϕ = f̂. Then (3.1) can be inferred from the equivalent [Co, Theorem 1], or shown as follows. Let n > 1. For each neighbor u of v, every vertex w ∈ C_{u,n} is at distance n + 1 or n − 1 from v, according if it is closer to u or to v. In the former case there is exactly one neighbor u of v for which w ∈ C_{u,n}, and this yields the first summand of the right-hand side. In the latter case there are q_{v} such neighbors u of v, namely all but the one that lies on the path from v to w; this accounts for the other summand. If n = 1, in this latter case w equals v, so it belongs to C_{u,1} for all the q_{v} + 1 neighbors u of v. Finally, if n = 0 the equality is immediate.

Equation (3.2) follows from the identity

C_{u,n} ∪ C_{v,n} = C_{uv,n} ∪ C_{uv,n−1} for uv ∈ E and n ∈ N

(without the last term if n = 0) and the fact that these are disjoint unions. Indeed, a vertex w ∈ C_{u,n} is, as before, at distance n + 1 or n − 1 from the neighbor v of u. In the former case it belongs to C_{uv,n}, in the latter to C_{uv,n−1}.

Finally, (3.2′) is verified as follows. Let n > 0. If u ≠ w ∼ v, then every vertex y ∈ C_{wv,n} is at distance n + 1, n, or n − 1 from uv, according if it is closest to w, v, or u. In the first case there is exactly one w as above, so we get the first summand in the bottom right-hand side. In the second case there are q_{v} − 1 such neighbors w of v, namely all except u and the one that lies on the path from v to y. In the third case there are q_{v} such neighbors w of v, namely all but u. If n = 0, in the second case y = v, so it belongs to C_{wv,0} for all the q_{v} neighbors w of v that are different from u.

Conversely, let ϕ satisfy (3.1), and set f(v) = ϕ(v, 0) for each v ∈ V. We want to prove by induction on n that f̂( · , n) = ϕ( · , n) on V. For n = 0 this holds by definition, so assume it proved up to n. Equation (3.1) has been proved to hold for f̂ in place of ϕ, and it expresses ϕ( · , n + 1) in terms of ϕ( · , n) and ϕ( · , n − 1). Since ϕ( · , k) = f̂( · , k) for k = n, n − 1, we conclude that this is also true for k = n + 1. If (3.2) also holds then we see that f̂( · , n) = ϕ( · , n) on E, again by induction in a similar fashion. Same if (3.2′) replaces (3.2).

The verifications concerning Z_{V} , Z_{E} are straightforward.

4. Circles through a fixed point

4.1. Inversion. The overdetermination described in Theorem 3.1 can be compensated by restricting the circle transform to a suitable subcomplex of the space C of all circles of V. One possibility is to take the circles that pass through a given reference vertex o. Since these can be parametrized by their centers, the resulting subcomplex is therefore in one-to-one correspondence with V ∪ E via x ↦ C_{x} = C_{x,|x|}. For functions f ∈ ℓ^{1}(V), the corresponding restriction R of the circle transform is given by

Rf(x) = f̂(x, |x|) = ∑_{u∈C_{x}} f(u) for every x ∈ V ∪ E.

The circles through o are not enough to separate all vertices of V. We shall describe a partition of V with the property that, if a vertex belongs to a circle, then so do all the elements of the same coset. One can therefore recover from Rf at most the sum of values of f at the vertices in each of these cosets. We shall actually prove that this is possible, and give the inversion formula.

On each given circle of positive radius through o, all vertices, except o itself, share the unit-length predecessor v_{1} of the center. Therefore it is enough to restrict attention to the subtree spanned by {o} ∪ S_{o′} for each fixed neighbor o′ of o. Thus in the sequel we shall assume without loss of generality that q_{o} = 0, that is, o has only one neighbor o′. In order to avoid exceptions in forthcoming formulas, set o_{1} = o′ (although o′ is not a predecessor of o) and S_{o′,−1} = {o}.

First consider the restriction R_{V} of R to v-circles (through o). Any such circle contains only even vertices, so R_{V} f depends only on the restriction of f to V_{even}. The cosets of V_{even} will be described as inverse images of single vertices by a sort of contraction map β, in terms of which R_{V} can be explicitly factored through an invertible transform.

We first describe v-circles more explicitly.

Lemma 4.1. If o ≠ v ∈ C_{x} for some x ∈ V, then v is even and x ∈ S_{v_{|v|/2}} \ S_{v_{|v|/2+1}}. Therefore, if a vertex u is such that

u_{|u|/2+1} = v_{|v|/2+1}

(in particular, |u| = |v|), then u ∈ C_{x}.

Proof. It is enough to realize that |x| ≥ |v|/2, and that the join of v, x is v_{|v|/2}.

Now we can give a natural definition of the contraction map β: V_{even} → V′ by

β(v) = v_{|v|/2+1} for every v ∈ V_{even}

(with the special case β(o) = o′). Since every level set of β is finite, there is an induced operator B that maps a function f on V_{even} (or V) to the function Bf on V′ given by

Bf(w) = ∑_{v∈β^{−1}(w)} f(v) for every w ∈ V′.

The map β is easily seen to be onto, then so is the operator B. Moreover B maps ℓ^{1}(V_{even}) (or ℓ^{1}(V)) into ℓ^{1}(V′) with unit operator norm, its kernel is the set of summable functions on V_{even} whose sum vanishes on each level set of β, namely on

β^{−1}(w) = S_{w,|w|−2} for w ∈ V′,

and its range is the set of summable functions on V′. A right inverse of B is the operator B^{−1} that maps a function g on V′ to the function B^{−1}g on V_{even} given by

B^{−1}g(v) = \frac{g(β(v))}{#β^{−1}(β(v))} for every v ∈ V_{even}.

Such a right inverse also has unit ℓ^{1} operator norm, and its range is the space B of summable functions on V_{even} that are constant on each level set of β.

Let us describe the images of circles through the contraction.

Lemma 4.2. If v ∈ V, the image β(C_{v}) of the circle C_{v} under β is

C′_{v} = {o′} ∪ {w ∈ V′: w ∉ [o, v], w^{−} ∈ [o, v]}

(with the special case C′_{o} = {o′}). Indeed β^{−1}(C′_{v}) = C_{v}.

Proof. If o ≠ u ∈ C_{v}, let v_{n} = u_{n} be the join of v, u. Then |v| = d(v, u) = d(v, v_{n}) + d(v_{n}, u) = (|v| − n) + (|u| − n), so n = |u|/2. If w = β(u), a predecessor of u, then w^{−} = u_{n} ∈ [o, v], while w ∉ [o, v].

Conversely, if o′ ≠ w ∈ C′_{v} and if u ∈ β^{−1}(w) then w^{−} is the join of v, w, so d(v, u) = d(v, w^{−}) + d(w^{−}, w) + d(w, u) = (|v| − |w| + 1) + 1 + (|w| − 2) = |v|.

By Lemma 4.1, each circle through o is a union of level sets of β, hence we can factor R_{V} = R′B, where R′ is the Radon transform that integrates summable functions on V′ along the sets C′_{v}. Namely, if g ∈ ℓ^{1}(V′), then

R′g(v) = ∑_{w∈C′_{v}} g(w) for every v ∈ V.

Notice that |R′g| ≤ ∥g∥_{V′}, the ℓ^{1} norm of g. (Both transforms R_{V} , R′ take values among functions on V.) Certainly R_{V} is not injective, since ker R_{V} contains ker B. We shall prove that this is the only obstruction to invertibility—that is, R′ is injective, and the inverse of R_{V} is only determined up to averages on level sets of β, each of which is finite. In other words, we shall show how to invert R_{V} on B.

Proposition 4.3. If g is a function on V′ then

D^{−}R′g(w) = D^{+}g(w) + \begin{cases}g(w) & if w = o, o′, \\ 0 & if o, o′ ≠ w ∈ V. \end{cases}

Therefore D^{−}R′ maps ℓ^{1}(V′) into ℓ^{1}(V) with operator norm not greater than 2, and any function g ∈ ℓ^{1}(V′) can be expressed in terms of ϕ = R′g by

g(v) = \begin{cases}D^{−}ϕ(o) & if v = o′, \\ − ∑_{w∈S_{v}} D^{−}ϕ(w) & if o′ ≠ v ∈ V′, \end{cases} (4.1)

where the series converges absolutely. In particular, D^{−}R′, R′ are one-to-one. Moreover ϕ satisfies the relation

∑_{w≠a} D^{−}ϕ(w) = 0, (4.2)

again with absolute convergence.

Proof. The first equality holds in the case w = o because C′_{o} = {o′}, whereas, if w ≠ o, because

C′_{w} \ C′_{w^{−}} = S_{w,1},
C′_{w^{−}} \ C′_{w} = \begin{cases}∅ & if w = o′, \\ \{w\} & if w ≠ o′. \end{cases}

The assertion about the norm is a consequence of the definition of D^{+} and, as in Lemma 2.1, of the uniqueness of the immediate predecessor u^{−1} of each given u ≠ o. As to (4.1), the case v = o′ is again immediate, whereas for o′ ≠ v ∈ V′ it follows from Lemma 2.1, as does relation (4.2).

Remark 4.4. The subtraction due to the difference operator D^{−} in the series in (4.1) cannot in general be distributed, namely as ∑_{w∈S_{v}} ϕ(w) − ∑_{w∈S_{v}} ϕ(w^{−}), to obtain cancelations.

An inversion for R_{V} can only work on a space, such as B, that has trivial intersection with its kernel. Such partial inversion can now be obtained by simply applying B^{−1} to (4.1).

Having inverted the restriction R_{V} of R to the set of v-circles (through o), let us now turn our attention to R itself. Lemma 4.1 carries on word by word if we replace x ∈ V with each of the endpoint vertices of x ∈ E. Therefore the partition on V_{even} described above, also in terms of the map β, is not properly refined by e-circles; in other words, these do not separate a pair of even vertices unless also v-circles do. Nonetheless, e-circles also contain odd vertices, which in turn divide into cosets. Each odd vertex v belongs to exactly one circle, namely the e-circle centered at the edge v_{(|v|+1)/2}v_{(|v|−1)/2}. Therefore V_{odd} gets partitioned into the cosets S_{w,|w|−1} for w ∈ V′. In order to get simple unified expressions, we define a map θ: V → V′ ∪ E so that the inverse images of vertices are these cosets, and that θ restricts to μ ∘ β on V_{even}. Indeed

θ(v) = \begin{cases}v_{|v|/2+1}v_{|v|/2} & if v ∈ V_{even}, \\ v_{(|v|+1)/2} & if v ∈ V_{odd}. \end{cases}

As with β, there is an induced operator Θ that maps a function f on V to the function Θf on V′ ∪ E given by

Θf(x) = ∑_{v∈θ^{−1}(x)} f(v) for every x ∈ V′ ∪ E.

Since θ is onto, then so is Θ, which maps ℓ^{1}(V) into ℓ^{1}(V′ ∪ E) with unit operator norm, its kernel is the set of summable functions on V whose sum vanishes on each level set of θ, namely on

θ^{−1}(x) = S_{x,|x|−1} for every x ∈ V′ ∪ E,

and its range is the set of summable functions on V′ ∪ E. A right inverse of Θ is the operator Θ^{−1} that maps a function h on V′ ∪ E to the function Θ^{−1}h on V given by

Θ^{−1}h(v) = \frac{h(θ(v))}{#θ^{−1}(θ(v))} for every v ∈ V.

Such a right inverse also has unit ℓ^{1} operator norm, and its range is the space of summable functions on V that are constant on each level set of θ.

The function Θf contains as much information on f as can be recovered from the knowledge of Rf. For f ∈ ℓ^{1}(V) and ϕ = Rf, we have shown in (4.1) that

Θf(x) = \begin{cases}D^{−}ϕ(o) & if x = o′o, \\ − ∑_{w∈S_{x}} D^{−}ϕ(w) & if o′o ≠ x ∈ E. \end{cases}

The sum of f on each coset in V_{odd} can also be recovered from Rf:

Proposition 4.5. If f is a function on V then

D^{−}Rf(x) = D^{+}Θf(x) + \begin{cases}Θf(x) & if x = o′o, \\ 0 & if o′o ≠ x ∈ E. \end{cases}

Therefore D^{−}R maps ℓ^{1}(V) into ℓ^{1}(V ∪ E) with operator norm not greater than 3. For any f ∈ ℓ^{1}(V) one can recover Θf on V′ in terms of ϕ = Rf by

Θf(v) = D^{−}ϕ(vv^{−}) − ∑_{w∈S_{v}} D^{−}ϕ(w) for v ∈ V′, (4.3)

where the series converges absolutely.

Proof. If v ∈ V′ then

C_{vv^{−}} \ C_{v^{−}} = S_{v,|v|−1} = θ^{−1}(v),
C_{v^{−}} \ C_{vv^{−}} = \begin{cases}∅ & if v = o′, \\ S_{v,|v|−2} = θ^{−1}(vv^{−}) & if v ≠ o′. \end{cases}

Therefore

D^{−}ϕ(vv^{−}) − Θf(v) = \begin{cases}0 & if v = o′, \\ −Θf(vv^{−}) & if v ≠ o′, \end{cases}

that equals ∑_{w∈S_{v}} D^{−}ϕ(w) in both cases, by (4.2) in the former, while applying (4.1) to g = Bf in the latter. As to norms, we have

∥D^{−}Rf∥_{V ∪E} = ∥D^{−}Rf∥_{V} + ∥D^{−}Rf∥_{E}
≤ 2∥f∥_{V_{even}} + (∥f∥_{V_{odd}} + ∥f∥_{V_{even}}).

This completes the description of the inversion of R.

4.2. Range. For every f ∈ ℓ^{1}(V), relation (4.2) is the only non-trivial one satisfied by Rf:

Proposition 4.6. The range of R (respectively R_{V}) on ℓ^{1}(V) is the space of functions ϕ on V ∪ E (respectively V) such that D^{−}ϕ is in ℓ^{1} of V ∪ E (respectively V), relation (4.2) holds, and the function g given by (4.1) is in ℓ^{1}(V′).

Proof. We first show that R′g = ϕ on V. Indeed, R′g(o) = g(o′) = D^{−}ϕ(o) = ϕ(o) because C′_{o} = {o′}, while if u ≠ o we have

R′g(u) = D^{−}ϕ(o) − ∑_{\begin{matrix}v∈C′_{u}\\{o′\} \\ w∈S_{v} \end{matrix}} D^{−}ϕ(w) = ∑_{w∈[o,u]} D^{−}ϕ(w) = ϕ(u)

by (4.2), since V is the disjoint union of [o, u] and of S_{v} for v ∈ C′_{u}\{o′}, and by (2.1). Then f = B^{−1}g ∈ ℓ^{1}(V_{even}) and Rf = R′g = ϕ on V.

Let h be the function on V′ given by

h(v) = D^{−}ϕ(vv^{−}) − \begin{cases}0 & if v = o′, \\ −g(v) & if v ≠ o′, \end{cases}

and extend to V_{odd} the f determined before by f = Θ^{−1}h. Then ∥f∥_{V_{odd}} = ∥h∥ ≤ ∥D^{−}ϕ∥ + ∥g∥, while Rf = ϕ on E is obtained by comparing the definition of h with (4.3), in view of (4.1), of the definition of D^{−}, and the relation Rf = ϕ already proved on V.

5. Some other subcomplexes

We conclude by observing that the restriction of the full circle transform to some other natural subcomplexes, inspired by the classical setting in the plane, is essentially equivalent to the situation described in Section 4.

One possible subcomplex is that of circles tangent to a geodesic K of V, that are those that intersect K in exactly one vertex each. This tangency point is also the closest vertex of K to the center as well as to each of the vertices of the circle. Therefore, one can restrict to the subtree of T spanned by the vertices of V whose closest vertex of K is a given vertex o ∈ K, and realize that on such subtree we have the situation of the preceding section. There follows a similar description of the lack of injectivity of the transform, a partial inversion, and a range characterization.

The same considerations carry on verbatim if we take circles externally tangent to any given connected subset K of V —if K is infinite, the adverb “externally” can be omitted. In order for every point of K to belong to some such circles one has to add the requirement that the interior of K be empty.

References

[BCCP] C. A. Berenstein, E. Casadio Tarabusi, J. M. Cohen, M. A. Picardello, Integral geometry on trees, Amer. J. Math. 113 (1991), 441–470. MR 92g:05066

[BCP] C. A. Berenstein, E. Casadio Tarabusi, M. A. Picardello, Radon transforms in hyperbolic spaces and their discrete counterparts, Telesio Courses in Math. (to appear).

[Ca] E. Casadio Tarabusi, Radon transforms on hyperbolic spaces and trees, Geometry Seminars 1996–1997 (Bologna), Univ. Stud. Bologna, Bologna, 1998, pp. 1–13. MR 99e:53107

[CCP] E. Casadio Tarabusi, J. M. Cohen, M. A. Picardello, The horocyclic Radon transform on non-homogeneous trees, Israel J. Math. 78 (1992), 363–380. MR 94b:44002

[Co] J. M. Cohen, Operator norms on free groups, Boll. Un. Mat. Ital. B (6) 1 (1982), 1055– 1065. MR 85d:22011

[J] F. John, The ultrahyperbolic differential equation with four independent variables, Duke Math J. 4 (1938), 300–322; reprinted in 75 years of Radon transform (Vienna, 1992), Conf. Proc. Lecture Notes Math. Phys., vol. 4, Internat. Press, Cambridge, 1994, pp. 301–323. Zbl 19.024, MR 96d:35096

[M] I. G. Macdonald, Spherical functions on a group of p-adic type, Publications of the Ramanujan Institute, vol. 2, Ramanujan Institute, Centre for Advanced Study in Mathematics, University of Madras, Madras, 1971. MR 55 #8261

[S] J.-P. Serre, Arbres, amalgames, SL_{2}, Astérisque, vol. 46, Soc. Math. France, Paris, 1977; English translation, Trees, Springer, Berlin, 1980. MR 57 #16426, 82c:20083

Dipartimento di Matematica “G. Castelnuovo”, Università di Roma “La Sapienza”, Piazzale A. Moro 2, 00185 Roma, Italy

  E-mail address: casadio@science.unitn.it

Department of Mathematics, Hill Center, Rutgers University, New Brunswick, NJ 08903, U.S.A.

  E-mail address: gindikin@math.rutgers.edu

Dipartimento di Matematica, Università di Roma “Tor Vergata”, Via della Ricerca Scientifica, 00133 Roma, Italy

  E-mail address: picard@mat.uniroma2.it